OpenStudy (anonymous):

How do I solve this? My algebra 2 teacher gave me the problem x+7/3 - 4x+1/9 and also gave us the answer, which is -x+20/9. I have missed a lot of school recently and I am totally lost. HELP!

OpenStudy (mathstudent55):

You need to combine like terms. Like terms are terms that have the same variable part. x and -4x are like terms and can be combined together. 7/3 and 1/9 are like terms and can be combined together.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

BUT that gives me 4x^2+8/11. Am I doing something wrong or is that just a step?

OpenStudy (mathstudent55):

x + 7/3 - 4x + 1/9 = = x - 4x + 7/3 + 1/9 First combine x with -4x. Then combine 7/3 with 1/9.

OpenStudy (anonymous):

x+7/3-4x+1/9=0 *9 9x+21-36x+1=0 -27x=-22 x=22/27

OpenStudy (mathstudent55):

When you add or subtract terms with x, the power of x does not change. For example: 2x + 5x - 9x = 7x - 9x = -2x

OpenStudy (mathstudent55):

@Genc How did you make of a smple collecting like terms problem into an equation?

OpenStudy (anonymous):

But I am not solving for x, I have to arrive at the answer that was provided.

OpenStudy (anonymous):

that was @Genc

OpenStudy (mathstudent55):

x + 7/3 - 4x + 1/9 = = x - 4x + 7/3 + 1/9 Do you see what I did above? I placed the terms with x near each other, and I did the same for the terms wiith no x. Now combine together x and -4x Then combine together 7/3 and 1/9

OpenStudy (anonymous):

ooooooooo

OpenStudy (anonymous):

I'm still not where I need to be.

OpenStudy (anonymous):

I'm really sorry this must be frustrating.

OpenStudy (anonymous):

@ConfusedSophia ,, do you have to find the X? in your problem ?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

x + 7/3 - 4x + 1/9 = = x - 4x + 7/3 + 1/9 Now combinbe x and -4x to get -3x Then combine 7/3 and 1/9. You need a common denominator of 9. 7/3 + 1/9 = 21/9 + 1/9 = 22/9 So x - 4x + 7/3 + 1/9 = -3x + 22/9 Your answer above is incorrect.

OpenStudy (anonymous):

hmmm ok. are you sure?

OpenStudy (anonymous):

Can I ask you about to help me with an inequality problem too, @mathstudent55 ?

OpenStudy (anonymous):

or @Genc too

OpenStudy (anonymous):

x+7/3 - 4x+1/9 = -3x+(7/3+1/9)= -3x+(22/9)= ........so if we divide with /3 we have = -x+22/27 ...this can be the correkt answer

OpenStudy (anonymous):

hmmm ok

OpenStudy (anonymous):

@mathstudent55 ...yeah it .does not matter anything..the result will be the same

OpenStudy (mathstudent55):

What? -3x + 22/9 is the same as -x + 22/9?

OpenStudy (anonymous):

@mathstudent55 -3x+22/9 is the same with -x+22/27

OpenStudy (mathstudent55):

If you really think it is, why don't you replace x with the same number in both and see if they are equal. How about try x = 0.

OpenStudy (mathstudent55):

(1/3)(-3x + 22/9) = -x + 22/27 -x + 22/27 is one third of -3x + 22/9. They are not equal.